# Form-preserving transformations of wave and Wigner functions

### v1

@@ 2 Remarkable time-dependent wave functions generated by form-preserving transformations
The Airy beam [1] and the Senitzky coherent states [27] are noteworthy solutions of the time-dependent Schrödinger equation
@@
Here $\psi(x,t)$ is the wave function and the general potential energy is $V=V(x,t)$.
@@
$\psi(x,t)\ =\ R\big(x-q(t)\big)\,e^{i\varphi(x,t)}\ ;\ \ R,\,\varphi,\,q(t)\,\in\,{\mathbb{R}}\ .$ (5)
We derive them both now by Senitzky’s method [27].
Substituting the ansatz (5) in (1) produces an equation with real part
$-\frac{\hbar^{2}}{2m}\partial_{x}^{\,2}R+\left[V+\hbar\,\partial_{t}\,\varphi\,+\,\frac{\hbar^{2}}{2m}\,\left(\partial_{x}\varphi\right)^{2}\right]R\ =\ 0\ ,$ (6)
and imaginary part
$-2\frac{m}{\hbar}\,\dot{q}\,\partial_{x}R\ +\ 2\,\partial_{x}\varphi\,\partial_{x}R\ +\partial_{x}^{2}\varphi\,R\ =\ 0\ .$ (7)
Here we write $\partial/\partial x=\partial_{x}$, etc. (note that $\partial_{x}R(x)=dR(x)/dx$), and an overdot denotes a total time derivative.
Eq. (7) is independent of the potential $V$ and ensures that the continuity equation for probability is obeyed. It can be rewritten as
$\partial_{x}\left[\left(-\frac{m}{\hbar}\,\dot{q}\,+\,\partial_{x}\varphi\right)R^{2}\right]\,=\,0\quad\Rightarrow\quad\left(-\frac{m}{\hbar}\,\dot{q}\,+\,\partial_{x}\varphi\right)R^{2}\,=\,f(t)\,$, (8)
for some $f(t)$. For a normalizable solution, $R\to 0$ as $x\to\pm\infty$ at all times, showing that $f(t)$ must vanish. The nonnormalizable Airy beam solution vanishes at one of $\pm\infty$ and the same conclusion applies. [^2] Any solution $R$ that is constrained to vanish at any given point $x$ for all time will force $f(t)$ to vanish. Therefore, in both cases, solving for $\varphi$ in (8) we get
$\varphi(x,t)=\frac{m}{\hbar}\,\dot{q}\,x+\frac{{{S(t)}}}{\hbar}\ ,$ (9)
for some real function S(t){{S}}(t). Substituting this into (6) results in
$-\frac{\hbar^{2}}{2m}\,\partial_{x}^{\,2}\,R\ +\ \,\left[V(x)\,+\,\frac{1}{2}\,m\,\dot{q}(t)^{\,2}\,+\,m\,{\ddot{q}(t)}\,x\,+\,\dot{{S}}(t)\right]\,R\ =\ 0\ .$ (10)
If this is a valid differential equation for $R\big(x-{q}(t)\big)$, then its coefficients must depend only on xq(t)=:Xx-{q}(t)=:X [27].
With $V=m\omega^{2}x^{2}/2$, the coefficient of $R$ in (10) becomes $m\omega^{2}X^{2}/2$ if q(t){q}(t) describes classical harmonic motion, $\ddot{q}=-\omega^{2}q$, and $\dot{{S}}+m{\dot{q}}^{2}/2-m\omega^{2}{q}^{2}/2$ equals a constant, E-E. Eq. (10) is then the time-independent Schrödinger equation for a harmonic oscillator with position $X$ and energy $E$. The quantized energies are $E=E_{n}=\hbar\omega(n+1/2)$, $n=0,1,2,\ldots$, and
$R(X)\ =\ \left(\frac{\sqrt{{m\omega/\pi\hbar}}}{2^{n}\,n!}\right)^{1/2}H_{n}\left(\sqrt{\frac{m\omega}{\hbar}}\,X\right)\,e^{-m\omega X^{2}/2\hbar}$ (11)
results. Including the phase, we obtain the wave functions of (5).
Senitzky’s method also leads to the Berry-Balazs solution. For $V=0$, the (nontrivial) coefficient of $R$ in (10) becomes
$m\,\ddot{q}\,X\,+\,\frac{1}{2}\,m\,\dot{q}^{\,2}\,+\,m\,{\ddot{q}}\,{q}\,+\,\dot{{S}}(t)\ .$ (12)
In the first term $\ddot{q}$ must be a constant (denote it $B^{3}/2m^{2}$), and the remaining terms must sum to another constant (call it E-E). Again, (10) becomes a $t$-independent Schrödinger equation, this time for a linear potential, with Airy-function solution:
$-\frac{\hbar^{2}}{2m}\frac{d^{2}R}{dX^{2}}\,+\,$ $\frac{B^{3}}{2m}\left(X-\frac{2mE}{B^{3}}\right)R\,=\,0$ (13)
$\Rightarrow\qquad$ $R(X)\,=\,\text{Ai}\left[\frac{B}{\hbar^{2/3}}\left(X-\frac{2mE}{B^{3}}\right)\right]\,.$ (14)
Putting $E=0$, we recover the Schrödinger wave function (2).
The appearance of the Airy function is intriguing. It recalls the case of a linear potential for which the stationary-state wave functions are Airy functions. Consider $x$ now to be a transformed coordinate, with $X$ the original, now denoted $x^{\prime}$.
For the Berry-Balazs solution one starts with a linear potential ${V^{\prime}}\propto{x^{\prime}}$ and its energy-eigenstate state Airy-function ($\times\,e^{-iRt/\hbar}$) solutions. The Airy beam is obtained by transforming to a uniformly accelerating frame, $x={x^{\prime}}+B^{3}t^{2}/2$, so that the linear potential transforms to $V=0$, and the force-free solution results.
@@
To be clear, we state here that we are assuming $x$ and $x^{\prime}$ are Cartesian coordinates, so the transformations (18) of interest are those that map the Schrödinger equation in Cartesian coordinates to another Schrödinger equation in Cartesian coordinates, with a possibly different potential. So a change of coordinates from Cartesian to non-Cartesian is not included. For example, $x^{\prime}=L\,\exp\left(x/L\right)$, with $L$ a fixed length scale, produces a differential equation with a linear term that is not present in Cartesian coordinates. Similarly, in $D=2$ space dimensions, switching to polar coordinates changes the so-called form. Such transformations are well understood; we exclude them to focus on the more interesting Cartesian-to-Cartesian form-preserving transformations that will generalize Galilean-relativistic transformations.
@@
As noted before, the Senitzky states propagate coherently in the harmonic potential, without dispersion. What is sometimes said of the coherent (ground) states, roughly, that the $x^{2}$ potential counteracts the dispersing nature of the $p^{2}$ kinetic energy, is true of Senitzky excited states as well.
@@ 3 Schrödinger form preservation in D space dimensions
Here $R_{ab}$ are the entries of an orthogonal rotation matrix $R$ (with determinant 1), so that
$R\,R^{T}\ =\ R^{T}R\ =\ I\ ,$ (40)
where $I$ denotes the $D\times D$ identity matrix. All of $R$, $\beta$ and $\gamma$ depend on $t$ only. We then have
@@
The term $R^{T}\dot{R}\,\vec{x}$ is problematic, however. For a matrix $G$, ${\vec{\nabla}}\left(\vec{x}^{T}G\,\vec{x}\right)\,=\,2\,G\,\vec{x}$. But $R^{T}\dot{R}$ is antisymmetric: applying $\frac{d}{dt}$ to (40), we obtain $R^{T}\dot{R}=-(R^{T}\dot{R})^{T}$. Hence
@@
we refer to it as the $\mathrm{U(1)}$ Schrödinger equation.
Transformations of the equation (53) were investigated by Takagi in [29], where it was argued that (39) describes the most general possible form-preserving transformation. [^3] One of us has also presented an argument for the generality of (39) in [8], where other results related to this work can be found. In [21] Nikitin independently classifies the symmetries of (53) for different specified vector and scalar potentials, and also investigates form-preserving transformations, calling them equivalence transformations. We will contribute here by describing the transformations in this case explicitly and generally.
@@
Inserting (45) into (55) gives
$\frac{\hbar^{2}}{2m}\left[\nabla^{2}\psi^{\prime}+\frac{2i}{\hbar}{\vec{\nabla}}\psi^{\prime}\cdot{\vec{\nabla}}\phi+\frac{i}{\hbar}\psi^{\prime}\nabla^{2}\phi-\psi^{\prime}({\vec{\nabla}}\phi/\hbar)^{2}\right]\\ +\frac{i\hbar}{m}\vec{A}\cdot({\vec{\nabla}}\psi^{\prime}+\frac{i}{\hbar^{2}}\psi^{\prime}\,{\vec{\nabla}}\phi)+\left(\frac{i\hbar{\vec{\nabla}}\cdot{\vec{A}}+\vec{A}^{2}}{2m}+V\right)\psi^{\prime}\\ =i\hbar\left(\frac{\dot{\sigma}}{\sigma}\psi^{\prime}+\psi^{\prime}\frac{i}{\hbar}\partialderivative{\phi}{t}+\partialderivative{\psi^{\prime}}{t}\right).\qquad\qquad$ (56)
As before, we must turn the derivatives into primed derivatives. Using (42, 43, 44) we find
@@
Comparing with the 1-dimensional result (20) we see both comparable and new kinds of terms. Unlike in the 1D case, this result will help determine both $\phi$ and the relation between $\vec{A}$ and $\vec{A}\,^{\prime}$. But we are not guaranteed that, for a given set of vector potentials and rotations, a scalar field $\phi(\vec{x},t)$ exists such that its gradient satisfies (58).
@@
But the potential function must be real, so the coefficient of the imaginary term on the right hand side must be zero since all of the functions here have been taken to be real. This is easily solved by setting
$\sigma=\gamma^{-\frac{D}{2}}\ .$ (68)
@@
Comparing the transformed potentials in the $D$-dimensional and 1-dimensional cases (71, 18), we see differences due to the vector potential $\vec{A}$ and the rotation $R$. Recall that the nonrelativistic Lagrangian for a particle in a magnetic field features an interaction potential $\vec{A}\cdot\vec{v}$ where $\vec{v}$ is the speed of the particle. The second-last term of the right-hand side of (71) is the interaction of the moving reference frame with the magnetic field. The last term is the kinetic energy of the rotation, as we will confirm in the next subsection.
@@ 3.1 Schrödinger form preservation in 3 dimensions
In this subsection we focus on the special case of $D=3$ spatial dimensions, and matrices $R$ in the $\mathrm{SO(3)}$ subgroup of $\mathrm{O(3)}$. To start, the $\dot{R}$ terms in the transformations (70, 71) will be rewritten in familiar $D=3$ vector-product notation.
@@
Compare this result with that obtained for $D=1$, equation (18). The top line is just the $D=3$ version of the $D=1$ result. The 1st of the 2 additional terms on the 2nd line is the kinetic energy of the rotating frame which creates the outwards centrifugal force. The second term is the interaction potential of the moving frame with the magnetic field, as discussed earlier.
@@ 4.1 Wigner function transformation from Schrödinger form preservation
Note that the Jacobian of the transformation $(x,p)\to({x^{\prime}},{p^{\prime}})$ is
${\left|\begin{matrix}\partial{x^{\prime}}/\partial x&\partial{x^{\prime}}/\partial p\cr\partial{p^{\prime}}/{\partial x}&\partial{p^{\prime}}/\partial p\end{matrix}\right|}={\left|\begin{matrix}1/\gamma&0\cr-m\dot{\gamma}&\gamma\end{matrix}\right|}=\ 1\ .$ (83)
(82) is therefore a linear canonical transformation:
$\left(\begin{matrix}x^{\prime}\cr p^{\prime}\end{matrix}\right)\ =\ \left(\begin{matrix}1/\gamma&0\cr-m\dot{\gamma}&\gamma\end{matrix}\right)\left(\begin{matrix}x\cr p\end{matrix}\right)\,+\,\left(\begin{matrix}\beta\cr m\gamma^{2}\dot{\beta}\end{matrix}\right)\ .$ (84)
@@ 4.2 Form-preserving transformations of the Moyal equation
Incidentally, the notation (108, 110) can be related to familiar expressions. With
@@
First, (127, 128) would follow from
@@
If ${t^{\prime}}={t^{\prime}}(t)$, then ${\partial}{x^{\prime}}/{\partial}x$ must be independent of $x$. That means $x^{\prime}$} is linear in $x$. We have been led back to transformations of the same type as (84).
@@ 5 Conclusion
Schrödinger form-preserving transformations were also studied in $D>1$. Confirming previous results [26, 29], we find that they work in a similar way as in $D=1$, when only a scalar potential is present. When time-dependent rotations are incorporated in the transformations, however, form preservation is possible only if a (transforming) vector potential is included [21]. Our treatment here focuses on determining the explicit, general form of the transformation (69-71) as was done for $D=1$ in (18) [12].
@@
Second, it seems surprising that the Schrödinger equation has a symmetry that is an extended time-dependent Galilean transformation with a time re-parametrization. Can one understand better the structure of the Schrödinger form-preserving transformations somehow? Is it a clue to the properties of the non-relativistic limit, with or without gravity? The recent review [11] makes clear that these and similar questions have been investigated for a long time now. Both group-theoretical/algebraic and geometric approaches (see also [9] and references therein) are being used.

### v2

@@ 2 Remarkable time-dependent wave functions generated by form-preserving transformations
Closed-form exact solutions of the time-dependent Schrödinger equation are harder to come by than those of the time-independent Schrödinger equation. This paper discusses a method of generating such solutions.
We first focus on a pair of striking wave functions.
2.1 Berry-Balazs and Senitzky solutions
The Airy beam [1] and the Senitzky coherent excited states [27] are noteworthy solutions of the time-dependent Schrödinger equation
@@ 2.1 Berry-Balazs and Senitzky solutions
Here $\psi(x,t)$ is the wave function and the potential energy is $V=V(x,t)$.
@@
$\psi(x,t)\ =\ R\big(x-q(t)\big)\,e^{i\varphi(x,t)}\ ;\ \ R,\,\varphi,\,q(t)\,\in\,{\mathbb{R}}\$, (5)
that will be explained below.
The appearance of the Airy function is intriguing. It recalls the case of a linear potential for which the stationary-state wave functions are Airy functions. Consider $x$ now to be a transformed coordinate, with the original coordinate denoted $x^{\prime}$.
For the Berry-Balazs solution one starts with a linear potential ${V^{\prime}}\propto{x^{\prime}}$ and its energy-eigenstate state Airy-function solutions. The Airy beam is obtained by transforming to a uniformly accelerating frame, $x={x^{\prime}}+B^{3}t^{2}/2$, so that the linear potential transforms to $V=0$, and the force-free solution results.
@@ 2.2 General form-preserving tranformations
We have seen that the Airy beam, the Senitzky coherent excited states, and freely dispersing harmonic waveforms can be understood as arising from form-preserving transformations of the time-dependent Schrödinger equation. From now on, our focus will be on the form-preserving transformations themselves rather than on particular solutions resulting from them.
@@ 3 Schrödinger form preservation in D space dimensions
Here $R_{ab}$ are the entries of an orthogonal rotation matrix $R$ (with determinant 1). All of $R$, $\beta$ and $\gamma$ depend on $t$ only. We then have
@@
The term $R^{T}\dot{R}\,\vec{x}$ is problematic, however. For a matrix $G$, ${\vec{\nabla}}\left(\vec{x}^{T}G\,\vec{x}\right)\,=\,2\,G\,\vec{x}$. But $R^{T}\dot{R}$ is antisymmetric: applying $\frac{d}{dt}$ to $R^{T}R=I$, we obtain $R^{T}\dot{R}=-(R^{T}\dot{R})^{T}$. Hence
@@
it can be called the $\mathrm{U(1)}$ Schrödinger equation.
Transformations of the equation (43) were investigated by Takagi in [29], where it was argued that (30) describes the most general possible form-preserving transformation. [^2] One of us has also presented an argument for the generality of (30) in [8], where other results related to this work can be found. In [21] Nikitin independently classifies the symmetries of (43) for different specified vector and scalar potentials, and also investigates form-preserving transformations, calling them equivalence transformations. We will contribute here by describing the transformations in this case explicitly and generally, in the manner of (9) [12]: see equations (58, 59, 60) below.
@@
Inserting (35) into (45) and using (32, 33, 34), we find
@@
Only the term involving $\gamma,\sigma$ is imaginary, and so with
$\sigma=\gamma^{-\frac{D}{2}}\$(57)
vanishes.
@@ 3.1 Schrödinger form preservation in 3 dimensions
In this subsection we focus on the special case of $D=3$ spatial dimensions, and matrices $R$ in the $\mathrm{SO(3)}$ subgroup of $\mathrm{O(3)}$. To start, the $\dot{R}$ terms in the transformations (59, 60) will be rewritten in familiar $D=3$ vector-product notation.
@@
Compare this result with that obtained for $D=1$, equation (9). The top line is just the $D=3$ version of the $D=1$ result. The 1st of the 2 additional terms on the 2nd line is the kinetic energy of the rotating frame which creates the outwards centrifugal force. The second term is the interaction potential of the moving frame with the magnetic field.
@@ 4.1 Wigner function transformation from Schrödinger form preservation
The Jacobian of this transformation is 1, so that (71) is a linear canonical transformation.
@@ 4.2 Form-preserving transformations of the Moyal equation
Incidentally, the notation (95, 97) can be related to familiar expressions. With
@@
First, (114, 115) would follow from
@@
If ${t^{\prime}}={t^{\prime}}(t)$, then ${\partial}{x^{\prime}}/{\partial}x$ must be independent of $x$. That means $x^{\prime}$} is linear in $x$. We have been led back to transformations of the same type as (71), i.e. linear and canonical.
@@ 5 Conclusion
Schrödinger form-preserving transformations were also studied in $D>1$. Confirming previous results [26, 29], we find that they work in a similar way as in $D=1$, when only a scalar potential is present. When time-dependent rotations are incorporated in the transformations, however, form preservation is possible only if a (transforming) vector potential is included [21]. Our treatment here focuses on determining the explicit, general form of the transformation (58-60) as was done for $D=1$ in (9) [12].
@@
Second, it seems surprising that the Schrödinger equation has a symmetry that is a time-dependent Galilean transformation extended by a certain type of time re-parametrization. Can one understand better the structure of the Schrödinger form-preserving transformations somehow? Is it a clue to the properties of the non-relativistic limit, with or without gravity? The recent review [11] makes clear that these and similar questions have been investigated for a long time now. Both group-theoretical/algebraic and geometric approaches (see also [9] and references therein) are being used.